# Shortcuts to adiabaticity for coherent atom transport in an adjustable family of two-dimensional optical lattices

### v1

@@ Abstract
Motivated by the compelling need for coherent atom transport in a variety of emerging quantum technologies, we investigate such transport on the example of an adjustable family of two-dimensional optical lattices [L. Tarruell et al., Nature (London) 483, 302 (2012)] that includes square, triangular, honeycomb, dimerized, and checkerboard lattices as its special cases; dynamical optical lattices of this type have already been utilized for the demonstration of topological pumping and the realization of two-qubit quantum gates with neutral atoms. At the outset, we propose the appropriate arrangements of acousto-optic modulators that give rise to a frequency imbalance between counterpropagating laser beams, thus leading to the dynamical-lattice effect in an arbitrary direction in the lattice plane. We subsequently obtain the dynamical-lattice trajectories that enable atom transport using shortcuts to adiabaticity (STA) in the form of inverse engineering based on a dynamical invariant of Lewis-Riesenfeld type. We then quantify the resulting atom dynamics using the transport fidelity computed from the numerical solutions of the relevant time-dependent Schrödinger equation. We do so for various choices of the system parameters and transport directions, finding favorable results for the achievable transport times and robustness of the resulting transport to various experimental imperfections.
@@ I Introduction
Recent developments in the realm of emerging quantum technologies – such as quantum sensing based on atom interferometry Nav ; D (a); Rod ; Prem ; Mart and neutral-atom quantum computing (QC) Jak ; Bri ; Wei (a); Moh ; Nem ; Kie ; Shi – as well as those in the area of quantum-state engineering Shi (2); Sto (a); Haa (1, 2); Nau ; Ste (1, 2); StoNau (1, 2) that facilitates those technologies, have underscored the need for a fast, nearly lossless, neutral-atom transport in optically-trapped ensembles Chia (c); Tor ; Che (b); Ste ; Zha ; Nes ; Lu+ ; Hic ; Din ; Hau ; HauSto ; Nor ; Hwa ; Mor ; Cic ; Ton . As a result, coherent atom transport Qic enabled by moving the confining optical trap – either an optical lattice, an array of optical tweezers, or their combination – has lately been attracting considerable attention, on the experimental- and theoretical sides alike. While the renewed experimental interest in this subject pertains largely to the issues related to increased scalability of neutral-atom QC platform Nor , the rekindled theoretical interest has also been motivated by the availability of powerful STA-based quantum-control protocols that can be utilized to model such transport STA ; STAf ; ContRev .
@@
The early theoretical investigations of coherent atom transport were mainly restricted to one-dimensional (1D) systems Tor with purely harmonic confining potentials Che (b), a physical situation that can be treated exactly within the LRI-based STA framework. However, given that such simplifications typically do not apply in realistic systems - where there is either a significant coupling between the longitudinal and transverse degrees of freedom, or the relevant confining potential is anharmonic Zha subsequent theoretical studies have been devoted to more practically relevant, higher-dimensional optical lattices of the optical-conveyor-belt Hau and double-well  HauSto types. Atom transport in such systems is enabled by the dynamical-lattice effect, which is induced by appropriately arranged pairs of acousto-optic modulators (AOMs) that give rise to a frequency imbalance between two counterpropagating laser beams. These studies of atom transport in higher-dimensional optical lattices Sto (b); Hof were carried out using the recently proposed eSTA method Whi (a); Odell (1, 2) - a predominantly analytical modification of the STA approach that allows one to account for the anharmonic character of the underlying potentials. The results obtained using this method are comparable to those corresponding to the quantum-speed limit HauSto .
@@
In this paper, we aim to demonstrate – using the STA approach – the feasibility of time-efficient coherent atom transport in an adjustable family of 2D optical lattices proposed in Tarr ; this family of lattices includes square, triangular, dimerized, honeycomb, and checkerboard lattices as its special cases. Dynamical optical lattices that belong to this family have already been utilized experimentally for demonstrating topological pumping Wal and for realizing two-qubit quantum gates with neutral atoms Kie ; these gates originate from exchange interactions between qubits Heu ; Heu (2); Sto (e); Hay .
@@
We start by proposing specific arrangements of AOMs that enable the dynamical-lattice effect in an arbitrary direction in the relevant lattice plane. We then set out to model the ensuing coherent atom transport by employing STA-type protocols, here based on inverse-engineering techniques involving LRIs. Having obtained the STA dynamical-lattice trajectories, we evaluate the resulting single-atom dynamics by solving the corresponding time-dependent Schrödinger equation (TDSE) in the comoving reference frame numerically using the Fourier split-operator method (FSOM). We then quantify the efficiency of STA-based atomic transport through the corresponding transport fidelity for a broad range of values of the relevant system parameters.
In addition to finding STA solutions to the atom-transport problem at hand, we devote significant attention to analyzing their sensitivity to errors, i.e. their robustness to various experimental imperfections, such as deviations in the control amplitudes, the duration of control, the unavoidable presence of noise, etc.
@@
The remainder of this paper is organized as follows. We start Sec. II by providing the basic background of the atom-transport problem at hand, including the Hamiltonian describing such transport and the essential properties of the relevant family of optical lattices; we then specify concrete arrangements of AOMs that enable atom transport in the system under consideration. Section III is devoted to a short survey of the general STA-related results of relevance for the problem at hand. In the following, Sec. IV, we describe how these general results are utilized for obtaining STA dynamical-lattice trajectories and analyze typical examples thereof. Section V is devoted to the description of our approach for the treatment of atomic dynamics. In Sec. VI, we discuss the obtained results for the atom-transport fidelities. To end with, a short summary of the paper - along with the main conclusions drawn - is provided in Sec. VII. In Appendix A we briefly review the application of the FSOM for solving TDSEs, while Appendix B contains a short introduction into the imaginary-time evolution (ITE) approach for computing ground states of quantum systems.
@@ II.1 Atom transport in optical lattices
Based on the Hamiltonian in Eq. (1), in Sec. IV below we discuss how – using an STA-based method – one can design trajectories that enable fast, coherent atom transport in the 2D optical-lattice system under consideration. In particular, both of our approaches (see Sec. II.3 below) enable atom transport along an arbitrary direction in the xyx-y plane; by way of example, we will mostly be interested in transport along the $x$-, $y$-, and diagonal directions, with the respective transport distances being denoted by $d_{\mathrm{x}}$, $d_{\mathrm{y}}$, and $d_{\mathrm{r}}$.
@@ II.2 Adjustable family of 2D optical lattices
In what follows, we describe the setup resulting in an adjustable family of 2D dynamical optical lattices that will be utilized as a test bed for the study of atom transport in the remainder of this work. In order to provide a more concise and clearer explanation of the envisioned setup, we hereafter restrict our discussion to the strong confining potentials in the xyx-y plane; we neglect the minor modulations arising from the ever-present confinement of the system in the $z$ direction, which in the anticipated physical realization is close to being harmonic.
@@
In the envisioned setup Tarr , beams $\overline{X}$ and $X$ propagate along the same axis, while beam $Y$ is oriented along a $90^{\circ}$-rotated axis (cf. Fig. 1). In particular, the interference of two perpendicular beams – namely, beams $X$ and $Y$ – creates a checkerboard lattice with period $\lambda_{\mathrm{L}}/\sqrt{2}$. At the same time, a third beam, $\overline{X}$ – which is collinear with $X$, but detuned from it by a frequency $\delta$ – gives rise to an additional standing wave with a period of $\lambda_{\mathrm{L}}/2$.
@@
- $2\alpha\,\sqrt{I_{X}I_{Y}}\sin(k_{\mathrm{L}}x)\cos(k_{\mathrm{L}}y)\cos\phi\>,$
where $\phi$ corresponds to the relative phase between the two arms, while $\alpha$ is the visibility introduced to quantify the influence of beam-quality in an experimental setting. The possible sources of experimental imperfections are, for instance, imperfect coherence and mode-mismatching. In what follows, we will restrict our analysis to high-contrast regimes of $\alpha=0.9$, which is readily achievable in experimental realizations Tarr .
@@
- $2\alpha\,\sqrt{U_{X}U_{Y}}\sin(k_{\mathrm{L}}x)\cos(k_{\mathrm{L}}y)\cos\phi\>,$
@@
The general optical-lattice potential in Eq. (II.2) gives rise – by adjusting the relative magnitude of the potential depths $U_{X}$, $U_{\overline{X}}$, and $U_{Y}$ – to several different lattice configurations; their corresponding potentials – as well as their respective periodically-repeating units (i.e. unit cells) – are visualized in Fig. 2. Among them, the following four lattice structures can be realized: square, dimerized, honeycomb, and 1D chains (cf. Fig. 2).
Figure 2: Density plots of the optical potential for the (a) square, (b) dimerized, (c) honeycomb, and (d) 1D-chains, optical lattices, obtainable by adjusting the different lattice depths present in the expression for $U_{\mathrm{L}}(x,y)$ [cf. Eq. (II.2)]. For each lattice structure, the corresponding unit cell is indicated by the dashed white line. The point $(x_{0},y_{0})=(0,0)$, around which the harmonic approximation of the full potential is developed, lies in the center od the unit cell.
@@
Given that the adjustable optical-lattice potential in Eq. (II.2) has minima around $(x_{0},y_{0})=(0,0)$ for all previously discussed configurations (cf. Fig. 2), we will approximate the full potential around the origin. We do so assuming that $k_{\mathrm{L}}|x|\ll 1$, as well as $k_{\mathrm{L}}|y|\ll 1$. Based on these assumptions, we straightforwardly obtain
@@
$V_{\mathrm{d},0}=U_{Y}+U_{\overline{X}}\cos^{2}\left(\frac{\theta}{2}\right)\>.$ (8)
The squared frequencies in Eq. (7) are given by
$\omega_{\mathrm{x}}^{2}$ $=$ $\frac{2U_{X}k_{\mathrm{L}}^{2}}{m}\left(\frac{U_{\overline{X}}}{U_{X}}\cos\theta-1\right)\>,$ (9)
$\omega_{\mathrm{y}}^{2}$ $=$ $\frac{2U_{Y}k_{\mathrm{L}}^{2}}{m}\>,$
while the expression for the factor $a_{\mathrm{x}}$, which has dimensions of acceleration, reads
$a_{\mathrm{x}}=\frac{U_{\overline{X}}k_{\mathrm{L}}}{m}\left(\sin\theta-2\alpha\>\frac{\sqrt{U_{X}U_{Y}}}{U_{\overline{X}}}\cos\phi\right)\>.$ (10)
@@
Another important property of the approximated potential $V_{\mathrm{L}}$ in Eq. (7), of interest for our further discussion, is that – by contrast to the total lattice potential, which is nonseparable in the xyx-y plane (i.e. it cannot be written as a sum of terms that depend on only on spatial coordinate)– this approximated potential is separable; this last property greatly simplifies the analysis of atom transport governed by this approximated potential.
@@ II.3 AOM configurations for atom transport
To facilitate further discussion of atom transport in the family of optical lattices under consideration, we first specify the relevant arrangements of AOMs that enable transport along $x$, $y$, and diagonal ($r$) directions.
To be able to appreciate the specific configurations of AOMs that give rise to the dynamical-lattice effect in the system at hand (see Fig. 3 below), it is helpful to first recall how the use of an AOM (also known as a Bragg cell) gives rise to the dynamical-lattice effect. AOMs lead to a shift of an incoming frequency by applying an ultrasonic frequency onto a crystal, which results in a density modulation along that crystal. Accordingly, the incoming laser light encounters the density modulations effectively playing the role of a grating and, consequently, leading to Bragg diffraction. Thus, the use of the AOM leads to a frequency shift that is equal to the frequency of the ultrasonic sound wave Hua . In an actual optical setup, the AOMs are connected to different channels of the direct-digital-synthesizer (DDS), which allow one to control the frequencies $f_{1}$ and $f_{2}$ applied to the crystals in the AOMs. While one AOM turns the frequency down, the other one ramps it upward by the same amount; this leads to an imbalance $\Delta f=f_{1}-f_{2}=c\left(|\mathbf{k}_{+}|-|\mathbf{k}_{-}|\right)/\left(2\pi\right)$ between the counterpropagating laser beams with wave-numbers $|\mathbf{k}_{+}|$ and $|\mathbf{k}_{-}|$, respectively. This imbalance engenders a dynamical (moving) optical lattice with velocity $v=\pi\Delta f/k_{\mathrm{L}}$, where $k_{\mathrm{L}}=|\mathbf{k}_{\mathrm{L}}|$ is the laser wave number HauSto .
@@
Having described the general procedure for incorporating AOMs into the system to enable atom transport, we now discuss how this approach can be implemented in the setup for the adjustable potential shown in Fig. 1. Currently, a single AOM is employed to tune the phase shift $\theta$ present in the optical-lattice potential of Eq. II.2. However, it is important to clarify that this particular AOM does not participate in the shifting procedure. To enable controlled lattice translation, each arm of the setup must first be modified so that the beams are not simply reflected by a mirror at the end of each arm. Instead, additional beam splitters and mirrors must be incorporated to separate the incoming and reflected components in each arm, allowing them to be treated independently. An illustration of the modified configuration for the relevant region is provided in Fig. 3. Once the four incoming beams are separated, individual AOMs can be placed in each path to control the frequency of the corresponding beam. In this way, a dynamical optical lattice can be generated in each arm through an induced frequency difference $\Delta f$. The resulting transport directions can then be represented as a superposition of displacements along each axis. For ideal diagonal transport, the frequency differences in both arms must be tuned identically.
III STA: survey of results relevant for atom transport
To set the stage for modeling coherent atom transport in the family of optical lattices under consideration using the STA- and eSTA methods (see Sec. IV below), we first provide a short survey of these methods, with emphasis on results of relevance for atom transport. We start with a short introduction into the subclass of STA methods that employ inverse engineering based on LRIs (Sec. III.1) and then specialize our discussion to the problem of harmonic transport (Sec. III.2); for the sake of simplicity, we provide an overview of these STA results using the notation that corresponds to 1D systems.

### v2

# Enhanced shortcuts to adiabaticity for coherent atom transport in a family of two-dimensional dynamical optical lattices

@@ Abstract
In view of the compelling need for coherent atom transport as a prerequisite for a variety of emerging quantum technologies, we investigate such transport on the example of an adjustable family of two-dimensional optical lattices [L. Tarruell et al., Nature (London) 483, 302 (2012)] that includes square, honeycomb, dimerized, and 1D-chains lattices as its special cases; dynamical optical lattices of this type have already been utilized for the demonstration of topological pumping and the realization of two-qubit quantum gates with neutral atoms. At the outset, we propose the appropriate arrangements of acousto-optic modulators that give rise to a frequency imbalance between counterpropagating laser beams, thus leading to the dynamical-lattice effect in an arbitrary direction in the lattice plane. We subsequently obtain the dynamical-lattice trajectories that enable atom transport in the lattices under consideration using two classes of control schemes: (i) shortcuts to adiabaticity (STA) in the form of inverse engineering based on a dynamical invariant of Lewis-Riesenfeld type, and (ii) their modification, known as enhanced STA (eSTA), which is well-suited for the treatment of anharmonic trapping potentials. We then quantify the resulting atom dynamics using transport fidelities computed from the numerical solutions of the relevant time-dependent Schrödinger equations. By doing so for various choices of the system parameters and transport directions, we demonstrate that – except in the special case of the dimerized lattice – the eSTA method consistently outperforms its STA counterpart, both in terms of the achievable transport times and the robustness of the resulting transport against small variations of optical-lattice depths.
@@ I Introduction
Recent developments in the realm of emerging quantum technologies – such as quantum sensing based on atom interferometry Nav ; D (a); Rod ; Prem ; Mart and neutral-atom quantum computing (QC) Jak ; Bri ; Wei (a); Moh ; Nem ; Kie ; Shi – as well as those in the area of quantum-state engineering Shi (2); Sto (a); Haa (1, 2); Nau ; Ste (1, 2); StoNau (1, 2) that facilitates those technologies, have underscored the need for a fast, nearly lossless, neutral-atom transport in optically-trapped ensembles Chia (c); Tor ; Che (b); Ste ; Zha ; Nes ; Lu+ ; Hic ; Din ; Hau ; HauSto ; Nor ; Hwa ; Mor ; Cic . As a result, coherent atom transport Qic enabled by moving the confining optical trap – either an optical lattice, an array of optical tweezers, or their combination – has lately been attracting considerable attention, on the experimental- and theoretical sides alike. While the renewed experimental interest in this subject pertains largely to the issues related to increased scalability of neutral-atom QC platform Nor , the rekindled theoretical interest has also been motivated by the availability of powerful STA-based quantum-control protocols that can be utilized to model such transport STA ; STA ; STA ; STAf ; ContRev .
@@
Early theoretical investigations of coherent atom transport were mainly restricted to one-dimensional (1D) systems Tor with purely harmonic confining potentials Che (b), a physical situation that can be treated exactly within the LRI-based STA framework. However, given that such simplifications typically do not apply in realistic systems - where either a significant coupling between the longitudinal and transverse degrees of freedom is present, or the relevant confining potential is anharmonic Zha - subsequent theoretical studies Hau ; HauSto have been devoted to more practically relevant, higher-dimensional optical lattices of the optical-conveyor-belt Hic and double-well Sto (b); Hof types. Atom transport in such systems is enabled by the dynamical-lattice effect, which is induced by appropriately arranged pairs of acousto-optic modulators (AOMs) that give rise to a frequency imbalance between two counterpropagating laser beams. These studies of atom transport in higher-dimensional optical lattices were carried out using the recently proposed eSTA method Whi (a); Odell (1, 2) - a predominantly analytical modification of the STA approach that allows one to account for the anharmonic character of the underlying potentials. The results obtained using this method are comparable to those corresponding to the quantum-speed limit HauSto .
@@
In this paper, we aim to demonstrate – using both the STA approach and eSTA-based modification – the feasibility of time-efficient coherent atom transport in an adjustable family of 2D optical lattices proposed in Tarr ; this family includes square, honeycomb, dimerized, and 1D-chains lattices as its special cases. Dynamical optical lattices that belong to this family have already been utilized experimentally for demonstrating topological pumping Wal and for realizing two-qubit quantum gates with neutral atoms Kie .
@@
We start by proposing specific arrangements of AOMs that enable the dynamical-lattice effect in an arbitrary direction in the relevant lattice plane. We then set out to model the ensuing coherent atom transport by employing two classes of control protocols: (i) STA-type protocols, here based on inverse-engineering techniques involving LRIs, and (ii) protocols based on the eSTA method, which unlike STA-based ones systematically take into account the full optical-lattice potential. Having obtained the dynamical-lattice trajectories based on each of these two methods, we evaluate the resulting single-atom dynamics by solving the corresponding time-dependent Schrödinger equation (TDSE) in the comoving reference frame numerically using the Fourier split-operator method (FSOM). We then quantify and compare the efficiency of STA- and eSTA-based atomic transport through the corresponding transport fidelities for a broad range of values of the relevant system parameters.
In the manner described above, we show that both approaches – STA and eSTA – allow time-efficient atom transport in the family of optical lattices under consideration. We demonstrate that while STA provides somewhat faster transport in shallow lattices (i.e. for relatively small lattice depths), eSTA becomes far superior to it for deeper lattices. Based on our numerical evaluations, we show that this superiority of the eSTA approach compared to its STA counterpart is more pronounced for shorter transport distances.
In addition to finding both STA- and eSTA solutions to the atom-transport problem at hand, we devote due attention to analyzing their sensitivity to errors, more precisely their robustness against deviations in optical-lattice depths. The obtained results bear out the superior robustness of the eSTA protocols compared to their STA counterparts in the transport problem under consideration, which to a large extent originates from the general assumptions underlying the inner workings of the eSTA approach.
@@
The remainder of this paper is organized as follows. We start Sec. II by providing the basic background of the atom-transport problem at hand, including the Hamiltonian describing such transport and the essential properties of the relevant family of optical lattices; we then specify concrete arrangements of AOMs that enable atom transport in the system under consideration. Section III is devoted to a short survey of the general STA- and eSTA-related results of relevance for the problem at hand. In the following, Sec. IV, we describe how these general results are utilized for obtaining STA and eSTA dynamical-lattice trajectories and analyze typical examples thereof. Section V is devoted to the description of our approach for the treatment of atomic dynamics. In Sec. VI, we discuss the results for the atom-transport fidelities obtained using either of the two approaches; we then discuss how these results are affected by the presence of small lattice-depth variations, and, finally, also demonstrate the superiority of our adopted transport schemes to some alternative approaches for modeling coherent atom transport in optical lattices. To end with, a short summary of the paper - along with the main conclusions drawn - is provided in Sec. VII. In order not to interrupt the flow of the discussion in the main part of this paper, certain cumbersome mathematical derivations are relegated to Appendices A and B. For similar reasons, in Appendix C we briefly review the application of the FSOM for solving TDSEs, while Appendix D contains a short introduction into the imaginary-time evolution (ITE) approach for computing ground states of quantum systems. Finally, in Appendix E we derive a dimensionless set of units that enables more accurate numerical simulations than the original units.
@@ II.1 Atom transport in optical lattices
Based on the Hamiltonian in Eq. (1), in Sec. IV below we discuss how – using STA- and eSTA methods – one can design trajectories that enable fast, coherent atom transport in the 2D optical-lattice system under consideration. In particular, both of our approaches (see Sec. II.3 below) enable atom transport along an arbitrary direction in the xyx-y plane; by way of example, we will mostly be interested in transport along the $x$-, $y$-, and diagonal directions, with the respective transport distances being denoted by $d_{\mathrm{x}}$, $d_{\mathrm{y}}$, and $d_{\mathrm{r}}$.
@@ II.2 Adjustable family of 2D optical lattices
In what follows, we describe the setup resulting in an adjustable family of 2D dynamical optical lattices that will be utilized as a test bed for the study of atom transport in the remaining part of this work. In order to provide a more concise and clearer explanation of the envisioned setup, we hereafter restrict our discussion to the strong confining potentials in the xyx-y plane; we neglect the minor modulations arising from the ever-present confinement of the system in the $z$ direction, which in the anticipated physical realization is close to being harmonic.
@@
In the envisioned setup Tarr , beams $\overline{X}$ and $X$ propagate along the same axis, while beam $Y$ is oriented along a $90^{\circ}$-rotated axis. In particular, the interference of two perpendicular beams – namely, beams $X$ and $Y$ – creates a checkerboard lattice with period $\lambda_{\mathrm{L}}/\sqrt{2}$. At the same time, a third beam, $\overline{X}$ – which is collinear with $X$, but detuned from it by a frequency $\delta$ – gives rise to an additional standing wave with a period of $\lambda_{\mathrm{L}}/2$.
@@
- $2\gamma\,\sqrt{I_{X}I_{Y}}\sin(k_{\mathrm{L}}x)\cos(k_{\mathrm{L}}y)\cos\phi\>,$
where $\phi$ corresponds to the relative phase between the two arms, while $\gamma$ is the visibility introduced to quantify the influence of beam-quality in an experimental setting. The possible sources of experimental imperfections are, for instance, imperfect coherence and mode-mismatching. In what follows, we will restrict our analysis to high-contrast regimes of $\gamma=0.9$, which is readily achievable in experimental realizations Tarr .
@@
- $2\gamma\,\sqrt{U_{X}U_{Y}}\sin(k_{\mathrm{L}}x)\cos(k_{\mathrm{L}}y)\cos\phi\>,$
@@
The general optical-lattice potential in Eq. (II.2) gives rise – by adjusting the relative magnitude of the potential depths $U_{X}$, $U_{\overline{X}}$, and $U_{Y}$ – to several different lattice configurations; their corresponding potentials – as well as their respective periodically-repeating units (i.e. unit cells) – are visualized in Fig. 2. Among them, the following four lattice structures can be realized: honeycomb, triangular, dimerized, and square. These four lattice types will be utilized in the following to test the performance of the eSTA scheme compared to its STA counterpart in modeling coherent atom transport.
Figure 2: Density plots of the optical potential for the (a) square, (b) dimerized, (c) honeycomb, and (d) 1D-chains, optical lattices, obtainable by adjusting the different lattice depths present in the expression for $U_{\mathrm{L}}(x,y)$ [cf. Eq. (II.2)]. For each lattice structure, the corresponding unit cell is indicated by the dashed white line. The point $(x_{0},0)$, around which the harmonic approximation of the full potential is developed, lies in the center of the unit cell.
@@
Given that the adjustable optical-lattice potential in Eq. (II.2) has a minimum at the point $(x_{0},y_{0})$ with
$x_{0}$ $=\frac{1}{k_{\mathrm{L}}}\arcsin\left(\gamma\frac{\sqrt{U_{X}U_{Y}}}{U_{\overline{X}}-U_{X}}\right)\>,$ (7)
$y_{0}$ $=0\>,$ (8)
for all previously discussed lattice configurations (see Fig. 2) we will approximate the full potential around their respective minima. We do so assuming that $k_{\mathrm{L}}|x-x_{0}|\ll 1$, as well as $k_{\mathrm{L}}|y|\ll 1$. Based on these last assumptions, it is straightforward to obtain
@@
$V_{\mathrm{d},0}=$ $U_{X}\sin^{2}(kx_{0})+\frac{U_{\overline{X}}}{2}\left[1+\cos(2kx_{0})\cos\theta\right]$ (10)
$+U_{Y}+2\gamma\sqrt{U_{X}U_{Y}}\sin(kx_{0})\cos\phi\>.$
The explicit expressions for the squared frequencies in Eq. (9) read
$\omega_{\mathrm{x}}^{2}$ $=$ $\frac{2U_{X}k_{\mathrm{L}}^{2}}{m}\left(\frac{U_{\overline{X}}}{U_{X}}\cos\theta-1\right)\cos^{2}(kx_{0})\>,$ (11)
$\omega_{\mathrm{y}}^{2}$ $=$ $\frac{2U_{Y}k_{\mathrm{L}}^{2}}{m}\left[1+\sqrt{\frac{U_{X}}{U_{Y}}}\>\gamma\sin(k_{\mathrm{L}}x_{0})\cos\phi\right]\>,$
while the factor $a_{\mathrm{x}}$, with dimensions of acceleration, is given by
$a_{\mathrm{x}}=-\omega_{\mathrm{x}}^{2}x_{0}\>.$ (12)
@@
Another important property of the approximated potential $V_{\mathrm{L}}$ in Eq. (9), of interest for our further discussion, is that – by contrast to the total lattice potential, which is nonseparable in the xyx-y plane (i.e. it cannot be written as a sum of terms that depend on only on spatial coordinate)– this approximated potential is separable; this last property greatly simplifies the analysis of atom transport governed by the approximated potential.
@@ II.3 AOM configurations for atom transport
To facilitate further discussion of atom transport in the family of optical lattices under consideration, we first specify the relevant arrangements of AOMs that enable transport along $x$ and $y$ directions (for an illustration, see Figs. 1 and 3).
@@
To be able to appreciate the proposed configurations of AOMs, it is helpful to first recall how the use of an AOM gives rise to the dynamical-lattice effect. AOMs lead to a shift an incoming frequency by applying an ultrasonic frequency onto a crystal, which results in a density modulation along that crystal. Accordingly, the incoming laser light encounters the density modulations effectively playing the role of a grating and, consequently, leading to diffraction. Thus, the use of the AOM leads to a frequency shift that is equal to the frequency of the ultrasonic sound wave Hua . In an actual optical setup, the AOMs are connected to different channels of the direct-digital-synthesizer (DDS), which allow one to control the frequencies $f_{1}$ and $f_{2}$ applied to the crystals in the AOMs. While one AOM turns the frequency down, the other one ramps it upward by the same amount; this leads to an imbalance $\Delta f=f_{1}-f_{2}=c\left(|\mathbf{k}_{+}|-|\mathbf{k}_{-}|\right)/\left(2\pi\right)$ between the counterpropagating laser beams with wavenumbers $|\mathbf{k}_{+}|$ and $|\mathbf{k}_{-}|$, respectively. This imbalance engenders a moving optical lattice with velocity $v=\pi\Delta f/k_{\mathrm{L}}$, where $k_{\mathrm{L}}=|\mathbf{k}_{\mathrm{L}}|$ is the laser wavenumber HauSto .
Having described the general procedure for incorporating AOMs into the system to enable atom transport, we now discuss how this approach can be implemented in the setup for the adjustable potential shown in Fig. 1. Currently, a single AOM is employed to tune the phase shift $\theta$ present in the optical-lattice potential of Eq. (II.2). However, it is important to clarify that this particular AOM does not participate in the shifting procedure. To enable controlled lattice translation, each arm of the setup must first be modified so that the beams are not simply reflected by a mirror at the end of each arm. Instead, additional beam splitters and mirrors must be incorporated to separate the incoming and reflected components in each arm, allowing them to be treated independently. An illustration of the modified configuration for the relevant region is provided in Fig. 3. Once the four incoming beams are separated, individual AOMs can be placed in each path to control the frequency of the corresponding beam. In this way, a dynamical optical lattice can be generated in each arm through an induced frequency difference $\Delta f$. The resulting transport directions can then be represented as a superposition of displacements along each axis. For ideal diagonal transport, the frequency differences in both arms must be tuned identically.
III STA and eSTA: survey of results relevant for atom transport
To set the stage for modeling coherent atom transport in the family of optical lattices under consideration using the STA- and eSTA methods (see Sec. IV below), we first provide a short survey of these methods, with emphasis on results of relevance for atom transport. We start with a short introduction into the subclass of STA methods that employ inverse engineering based on LRIs (Sec. III.1) and then specialize our discussion to the problem of harmonic transport (Sec. III.2); for the sake of simplicity, we provide an overview of these STA results using the notation that corresponds to 1D systems. Finally, we discuss the basic aspects of the eSTA method, explaining in detail how an eSTA control scheme can be obtained from its STA counterpart that is exact for a simplified Hamiltonian of the system (Sec. III.3).